# SkillWrapper: Generative Predicate Invention for Skill Abstraction

### v3

@@ Abstract
Generalizing from individual skill executions to solving long-horizon tasks remains a core challenge in building autonomous agents. A promising direction is learning high-level, symbolic representations of the low-level skills of the agents, enabling reasoning and planning independent of the low-level state space. Recent advances in foundation models have made it possible to generate symbolic predicates that operate on raw sensory inputs—a process we call generative predicate invention—to facilitate downstream representation learning. However, it remains unclear which formal properties the learned representations must satisfy, and how they can be learned to guarantee these properties. In this paper, we address both questions by presenting a formal theory of generative predicate invention for skill abstraction, resulting in symbolic operators that can be used for provably sound and complete planning. Within this framework, we propose SkillWrapper, a method that leverages foundation models to actively collect robot data and learn human-interpretable, plannable representations of black-box skills, using only RGB image observations. Our extensive empirical evaluation in simulation and on real robots shows that SkillWrapper learns abstract representations that enable solving unseen, long-horizon tasks in the real world with black-box skills.
1 Introduction
@@
An autonomous agent operating in the real world must process low-level sensory and motor signals while reasoning about high-level objectives (Doncieux et al., 2018; Konidaris, 2019). Analogous to how humans can perform complex tasks, like cooking or cleaning, without reasoning about muscle-level control, agents should have internal models of their skills that abstract away nuanced activities on the lower level. Such models must capture the necessary conditions for a skill to be executed (e.g., “pouring a teapot requires holding it first”) and the consequences of doing so (e.g., “pouring a teapot leaves it empty”). These two properties, known as preconditions and effects in the AI planning literature, enable compositional reasoning to identify long-horizon plans that can sequence lower-level skills to solve a task. Typically, these models must be specified manually. However, in real-world settings such skill representations may be nontrivial to acquire due to complex inter-skill constraints specific to the agent’s embodiment. This calls for algorithms that learn symbolic transition models of black-box skills without hand specification, enabling agents to directly utilize those skills to solve long-horizon tasks with off-the-shelf AI planners.
@@
Traditional approaches of skill abstraction often require factorizing the low-level state space to learn classifiers for each symbolic representation, relying heavily on hand-collected transition data (Konidaris et al., 2018). Recently, foundation models have enabled a new paradigm: generating semantically meaningful predicates directly from raw observations and directly evaluating their truth values on low-level observations (e.g., RGB images)—a process we refer to as generative predicate invention. Recent work has explored how foundation models can be used for predicate invention, by generating Python code to implement predicates (Liang et al., 2025) or sampling large predicate pools followed by sub-selection. However, these methods produce ad-hoc planning representations that cannot be guaranteed to solve a given task, and leave core questions on predicate invention unanswered: what properties should these learned abstractions satisfy, and how can they be learned to achieve these properties?
@@
Our answer is twofold. First, we develop a formal theory of generative predicate invention for skill abstraction, precisely characterizing the conditions under which a learned skill representation will be provably sound and complete with respect to downstream planning. Building on this foundation, we introduce SkillWrapper, a method explicitly designed to guarantee these theoretical criteria. SkillWrapper uses foundation models in three ways: interactively collecting data in the environment, proposing predicates when the current model fails, and classifying predicate truth values based solely on RGB image observations. Using these data and predicates, SkillWrapper learns symbolic representations of black-box skills that are both human-interpretable and directly usable for AI planning.
@@
We highlight the following contributions: (1) A formal theory of generative predicate invention for provably sound and complete skill abstraction; (2) SkillWrapper, a principled system built on this framework that leverages foundation models to learn interpretable symbolic representations of black-box skills; and (3) an extensive empirical evaluation of the system, demonstrating effectiveness in simulation and on two real robots.
@@
Figure 1: Overview of  SkillWrapper. For an agent equipped with black-box skills, SkillWrapper learns skill representations that are compatible with off-the-shelf planners. These representations are comprised of predicates invented by the foundation model. Given a novel planning problem described using the initial state and goal state as RGB images, a foundation model produces the corresponding abstract states by applying the invented predicates to the low-level states. SkillWrapper is agnostic to the agent, and we illustrate both real-world (robots) and simulated agents in this figure.
2 Problem Setting
In this section, we briefly discuss our problem setting while defining it formally in Appendix A. We consider an agent equipped with a finite set of object-centric skills $\Omega$, modeled as black-box options (Sutton et al., 1999). The agent can execute any $\omega\in\Omega$ and determine whether it succeeds, but it does not possess a symbolic transition model of these skills. Without such a model, the agent cannot plan over a long horizon without reasoning at the low-level state space $\mathcal{S}$}, which is continuous, high-dimensional, and impractical for classical search. The goal of Skill Model Learning is to acquire a symbolic abstraction of skills that enables efficient composition via classical planning.
Environment.
An environment is a tuple $({\mathcal{S}},{\mathcal{T}},\Omega,T)$, where $\mathcal{S}$} is the continuous state space, $\mathcal{T}$} is a finite set of object types, and $\Omega$ is the skill library. Each skill $\omega\in\Omega$ is parameterized by object types drawn from the set $\mathcal{T}$}. The environment dynamics are governed by an unknown transition function $T:{\mathcal{S}}\times\Omega\to{\mathcal{S}}$. A setting is defined as $(s_{0},{\mathcal{O}})$, consisting of an initial state $s_{0}\in{\mathcal{S}}$ and a set of typed objects $\mathcal{O}$} with $\tau(o)\subseteq{\mathcal{T}}$ for each $o\in{\mathcal{O}}$.
Black-box skills.
A skill $\omega\in\Omega$ is a tuple $(\mathcal{I}_{\omega},\pi_{\omega},\beta_{\omega},\theta_{\omega})$, where $\mathcal{I}_{\omega}\subseteq{\mathcal{S}}$ is the initiation set, $\pi_{\omega}$ is the option policy, $\beta_{\omega}\subseteq{\mathcal{S}}$ is the termination set, and $\theta_{\omega}=(\tau^{1}_{\omega},\ldots,\tau^{k}_{\omega})$ are the type constraints on its $k$ parameters. A skill instance $\underline{\omega}=\omega(o_{1},\ldots,o_{k})$ is valid for $o_{i}\in{\mathcal{O}}$ if $\tau^{i}_{\omega}\subseteq\tau(o_{i})$ for all $i$. Executing $\omega$ from $s\in\mathcal{I}_{\omega}$ terminates in some $s^{\prime}\in\beta_{\omega}$, affecting only the state of the bound objects.
Symbolic predicates and operators.
To enable abstract reasoning, we introduce a finite set of predicates $\mathcal{P}$}. Each $\sigma\in{\mathcal{P}}$ is parameterized by $\mathcal{T}$} and has a classifier $\phi_{\sigma}:{\mathcal{O}}^{n}\times{\mathcal{S}}\rightharpoonup\{0,1\}$ that evaluates its truth value for arguments in a state. Grounding predicates with valid objects induces a set of grounded predicates $\bar{\mathcal{P}}$, and hence an abstract state space $\bar{\mathcal{S}}=2^{\bar{\mathcal{P}}}$. The abstraction function $\Gamma:{\mathcal{S}}\to\bar{\mathcal{S}}$ maps each state to the set of grounded predicates true in that state, while the grounding function ${\mathcal{G}}:\bar{\mathcal{S}}\to{\mathcal{S}}$ maps an abstract state to the states in which its grounded predicates hold.
An operator $a\in{\mathcal{A}}$ is defined as $(\omega_{a},\theta_{a},\textsc{Pre}_{a},\textsc{Eff}_{a})$, where $\omega_{a}\in\Omega$ is the associated skill, $\theta_{a}$ are typed operator parameters, $\textsc{Pre}_{a}\subseteq{\mathcal{P}}$ is the precondition set, and $\textsc{Eff}_{a}=(\textsc{Eff}^{+}_{a},\textsc{Eff}^{-}_{a})$ are the add and delete effects. Grounding all operators $a\in{\mathcal{A}}$ with objects $o\in{\mathcal{O}}$ yields a set of abstract actions $\bar{a}\in\bar{\mathcal{A}}$, each executable whenever its ground preconditions hold in the current abstract state.
Planning problems.
Given an environment $({\mathcal{S}},{\mathcal{T}},\Omega,T)$, the agent’s objective is to learn an abstract transition model ${\mathcal{M}}=({\mathcal{P}},{\mathcal{A}})$, such that $\mathcal{M}$} is sufficient for planning with $\Omega$. Formally, a skill planning problem is $(s_{0},{\mathcal{O}},{\mathcal{S}}_{g})$, where $(s_{0},{\mathcal{O}})$ is a setting and ${\mathcal{S}}_{g}\subseteq{\mathcal{S}}$ are goal states. The corresponding abstract planning problem is $(\Gamma(s_{0}),\bar{\mathcal{A}},\bar{\mathcal{S}}_{g})$, where $\bar{\mathcal{S}}_{g}=\{\bar{s}\in\bar{\mathcal{S}}\mid\mathcal{G}({\bar{s}})\cap{\mathcal{S}}_{g}\neq\emptyset\}$. An abstract plan $\bar{\pi}=\langle\bar{a}_{1},\ldots,\bar{a}_{n}\rangle$ is valid if its execution under $\mathcal{M}$} yields an abstract trajectory consistent with some feasible low-level trajectory under $T$, ending in a goal state.
Problem statement.
The Skill Model Learning problem is: given experience in the form of stateskill-instancenext-state tuples $\langle s,\underline{\omega},s^{\prime}\rangle$, learn a model ${\mathcal{M}}=({\mathcal{P}},{\mathcal{A}})$ such that every plan found by a complete symbolic planner over $\mathcal{M}$} corresponds to a feasible low-level skill plan.
3 SkillWrapper
@@
In this section, we introduce SkillWrapper, a novel approach that autonomously learns symbolic representations for black-box skills using the concepts defined in Appendix C. To produce a valid abstract model that enables planning,  SkillWrapper iterates through a three-step process: 1) actively proposing and executing exploratory skill sequences to collect data on the initiation and termination set of each skill, 2) incrementally building a set of predicates from scratch by contrasting positive and negative examples, and then 3) constructing valid operators using these invented predicates, from which further exploratory skill sequences can be proposed. This procedure is outlined in Algorithm 1. As SkillWrapper continues to collect data, add predicates, and update its planning model, it learns a progressively more accurate abstract transition model. The resulting skill representations, or operators, can be used with an off-the-shelf classical planner to solve task planning problems specified using RGB images. We delve into the core components of our system (lines 4–9 of Algorithm 1) in the following subsections. Lastly, we provide strong theoretical results for the soundness and completeness of SkillWrapper in Section 3.4.
@@
9:end for
10:return $\mathcal{M}$
3.1 Active Data Collection
To collect data for learning, our method commands the agent to execute its skills in the world and then collects the resulting transitions. Each command is a sequence of skill instances $\langle\underline{\omega}_{1},\dots,\underline{\omega}_{k}\rangle$. While executing these commands, the agent collects a dataset $\mathcal{D}$ of transitions of the form $\langle s,\underline{\omega},s^{\prime}\rangle$, where $s,s^{\prime}\in{\mathcal{S}}$. These transitions can answer two questions:

### v4

# SkillWrapper: Generative Predicate Invention for Task-level Planning

@@ Abstract
Generalizing from individual skill executions to solving long-horizon tasks remains a core challenge in building autonomous robots. A promising direction is learning high-level, symbolic representations of the low-level skills of the robots, enabling reasoning and planning independent of the low-level state space. Recent advances in foundation models have made it possible to generate symbolic predicates that operate on raw sensory inputs—a process we call generative predicate invention—to facilitate downstream representation learning. However, it remains unclear which formal properties the learned representations must satisfy, and how they can be learned to guarantee these properties. In this paper, we address both questions by presenting a formal theory of generative predicate invention for task-level planning, resulting in symbolic operators useful for provably sound and complete planning. We propose SkillWrapper, a method that leverages foundation models to actively collect robot data and learn human-interpretable, plannable representations of black-box skills, using only RGB image observations. Our extensive empirical evaluation in simulation and on real robots shows that SkillWrapper learns abstract representations that enable solving unseen, long-horizon tasks in the real world with black-box skills.* Equal contribution. Work done primarily when the author was a postdoc researcher at Brown University.
I Introduction
@@
An autonomous robot operating in the real world must process low-level sensory and motor signals while reasoning about high-level objectives [12, 33]. Analogous to how humans can perform complex tasks, like cooking or cleaning, without reasoning about muscle-level control, robots should have internal models of their skills that abstract away from the nuances of low-level control. Such models must capture the necessary conditions for executing a skill (e.g., “pouring from a teapot requires holding it first”) and the consequences of doing so (e.g., “pouring from a teapot makes it empty”). These two properties, known as preconditions and effects in the AI planning literature [18], enable compositional reasoning to identify long-horizon plans that can sequence lower-level skills to solve a task. Typically, these models must be specified manually. However, in real-world settings, such task-level models may be nontrivial to acquire due to complex inter-skill constraints specific to the robot’s embodiment. This calls for algorithms that autonomously learn symbolic transition models of black-box skills, enabling robots to solve long-horizon tasks using off-the-shelf AI planners.
@@
Traditional approaches for learning task-level models often require either factorizing the low-level state space to learn classifiers for each symbolic representation [35], privileged information such as object poses [53], or initial pool of task-level symbols [57] (detailed in Appendix -F). Foundation models have enabled a new paradigm: generating semantically meaningful predicates from raw observations and directly evaluating their truth values on low-level observations (e.g., RGB images)—a process we refer to as generative predicate invention. While recent work has explored how foundation models can be used for predicate invention by generating Python code to implement predicates [39] or undirected sampling followed by a hill-climbing model search [4], these methods do not provide any concrete guarantees of correctness, leaving core questions unanswered: what properties should these learned abstractions satisfy, and how can they be learned to achieve these properties?
@@
Our answer is twofold. First, we develop a formal theory of generative predicate invention for task-level planning, precisely characterizing the conditions under which a learned model is provably sound and complete with respect to downstream planning. Building on this foundation, we introduce SkillWrapper, a method explicitly designed to guarantee these theoretical criteria. SkillWrapper uses foundation models in three ways: interactively collecting data in the robot’s environment, proposing predicates when the current model fails, and classifying predicate truth values based solely on RGB image observations. Using these data and predicates, SkillWrapper learns symbolic representations of black-box skills that are both human-interpretable and directly usable for task-level planning. To the best of our knowledge, SkillWrapper is the first work integrating active data collection and generative predicate invention using foundation models while provably guaranteeing soundness and completeness for task-level planning.
@@
We highlight the following contributions: (1) A formal theory of generative predicate invention for provably sound and complete representations; (2) SkillWrapper, a principled system built on this framework that leverages foundation models to learn interpretable symbolic representations of black-box skills; and (3) an extensive empirical evaluation of the system in simulated and real robotic platforms.
@@
Figure 1: Overview of  SkillWrapper. For a robot equipped with black-box skills, SkillWrapper learns task-level plans that are compatible with off-the-shelf planners. These representations are comprised of predicates invented by the foundation model. Given a novel planning problem described using the initial state and goal state as RGB images, a foundation model produces the corresponding abstract states by applying the invented predicates to the low-level states. SkillWrapper is agnostic to the robot embodiment, and we illustrate both real-world robots and simulation in this figure.
II Formal Framework
We consider a problem setting in which a robot is equipped with a set of predefined, “black-box” skills. The robot can evaluate whether a skill is executable in the current state, but it does not have a complete transition model of the skills a priori, and therefore cannot compose its skills to solve long-horizon problems without considering low-level details. However, if the robot were to learn a symbolic model of its skills, it could use classical task planning approaches to efficiently compose its skills and enable the robot to solve novel tasks. We formalize this setting as the problem of Skill Model Learning.
II-A Preliminaries
Environment model
We define an environment as a tuple $({\mathcal{S},\mathcal{T},\Omega,T})$, where the state space $\mathcal{S}$ is assumed to be high-dimensional, continuous, and fully observable; the type set $\mathcal{T}$ enumerates the possible object types; and the skills $\Omega$ are object-centric, such that each $\omega\in\Omega$ is parameterized by object types drawn from the type set $\mathcal{T}$. The transition function $T:\mathcal{S}\times\Omega\to\mathcal{S}$ characterizes the environment dynamics but is unknown to the robot. For example, Pour(?teapot, ?mug) may be used to pour tea from a teapot into a mug. However, certain environmental aspects (e.g., which mugs are available to pour into) may differ between settings during learning. We therefore define a specific setting by the tuple $({{s}_{0},\mathcal{O}})$, specifying an initial state ${s}_{0}\in\mathcal{S}$ and a set of typed objects $\mathcal{O}$, where the type(s) of object $o\in\mathcal{O}$ are denoted by $\texttt{t}_{o}\subseteq\mathcal{T}$.
Black-box skills
We model skills as object-centric options [60] with discrete, object-typed parameters. Formally, a skill $\omega\in\Omega$ is defined by a tuple $({\mathcal{I}_{\omega},\pi_{\omega},\beta_{\omega},\Theta_{\omega}})$, where the initiation set $\mathcal{I}_{\omega}\subseteq\mathcal{S}$ contains states from which the skill may be executed; the policy $\pi_{\omega}$ controls the robot during the skill; the termination set $\beta_{\omega}\subseteq\mathcal{S}$ is the set of states at which the skill immediately terminates; and the skill parameters $\Theta_{\omega}=({\theta_{\omega}^{1},\dots,\theta_{\omega}^{k}})$ specify type constraints $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\mathcal{T}$ for valid skill arguments. Specifically, a skill $\omega\in\Omega$ may be instantiated to create a skill instance $\underline{\omega}$ using objects ${o}_{1:{k}}\in\mathcal{O}^{k}$ if and only if for $1\leq i\leq k$, $\texttt{t}_{\theta_{\omega}^{i}}\subseteq\texttt{t}_{o_{i}}$. We assume that executing a skill only affects the state of objects passed as arguments. [^1] We do not, however, assume that a skill necessarily affects the state of all objects passed in as arguments.
Symbolic abstractions
Because the robot must evaluate $\mathcal{I}_{\omega}$ on individual states, it cannot distinguish essential skill information from irrelevant details, making long-horizon planning combinatorially difficult. However, each skill affects only a few objects at once, leaving the rest of the world unchanged. The robot can exploit this property by instead learning a factored state representation—formalized here as first-order logic models in PDDL [43]—providing an abstract transition model of its skills. Such abstractions enable the robot to use classical task planning to compose its skills and accomplish unseen, long-horizon goals.
We use symbolic predicates $\mathcal{P}$ to express abstract relations between objects. Each predicate $\sigma\in\mathcal{P}$ is a tuple $({C_{\sigma},\Theta_{\sigma}})$, where the predicate classifier $C_{\sigma}$ tests whether the predicate holds in a state, given a binding of objects to the predicate parameters $\Theta_{\sigma}=({\theta_{\sigma}^{1},\dots,\theta_{\sigma}^{n}})$. Each parameter $\theta_{\sigma}^{i}$ specifies a type constraint $\texttt{t}_{\theta_{\sigma}^{i}}\subseteq\mathcal{T}$ on corresponding object arguments. Formally, the partial function $C_{\sigma}:\mathcal{O}^{n}\rightharpoonup(\mathcal{S}\to\{0,1\})$ is defined on objects ${o}_{1:{n}}\in\mathcal{O}^{n}$ if and only if $\texttt{t}_{\theta_{\sigma}^{i}}\subseteq\texttt{t}_{o_{i}}$ for $1\leq i\leq n$. In such cases, we say that the objects ${o}_{1:{n}}$ are valid arguments for $\sigma$.
By observing different effects, the robot must construct a set of operators $\mathcal{A}$ that define an abstract transition model for the robot’s skills. We define each operator $a\in\mathcal{A}$ by $({\omega_{a},\Theta_{a},\textsc{Pre},\textsc{Eff}^{+},\textsc{Eff}^{-}})$, where $\omega_{a}\in\Omega$ is the corresponding skill; the operator parameters $\Theta_{a}=({\theta_{a}^{1},\dots,\theta_{a}^{m}})$ impose type constraints $\texttt{t}_{\theta_{a}^{i}}\subseteq\mathcal{T}$ on operator arguments; the preconditions Pre are a conjunction of literals over $\mathcal{P}$ defining the conditions necessary to apply the operator; and the add and delete effects, $\textsc{Eff}^{+}$ and $\textsc{Eff}^{-}$, are the subsets of $\mathcal{P}$ that become true and false, respectively, after the operator is applied.
Grounded Abstractions
To apply an abstract transition model to a real-world setting, the robot must map the low-level state to a high-level abstract state, which requires grounding the known predicates using concrete objects so that their truth value may be determined. A predicate $\sigma\in\mathcal{P}$ may only be grounded using valid arguments, inducing a grounded predicate $\underline{\sigma}=\sigma({o}_{1:{n}})$. Given a low-level state, the classifier $C_{\underline{\sigma}}:\mathcal{S}\to\{0,1\}$ tests whether the predicate holds for the object arguments. We define the set of grounded predicates for a setting $({{s}_{0},\mathcal{O}})$ under predicates $\mathcal{P}$ as $\underline{\mathcal{P}}=\{\sigma({o}_{1:{n}})\,:\,\sigma\in\mathcal{P},{o}_{1:{n}}\in\mathcal{O}^{n},\bigwedge_{i=1}^{n}\texttt{t}_{\theta_{\sigma}^{i}}\subseteq\texttt{t}_{o_{i}}\}$, inducing an abstract state space $\underline{\mathcal{S}}=2^{\underline{\mathcal{P}}}$ where each abstract state corresponds to a specific set of grounded relations.
II-B Problem Definition
We now define the task-level model learning problem and a task-level solution as follows:
Definition 1.
Given an environment $({\mathcal{S},\mathcal{T},\Omega,T})$ containing settings $\{({{s}_{0},\mathcal{O}})\}_{i=1}^{N}$, we define a Task-Level Model Learning problem as learning an abstract transition model $\mathcal{M}=({\mathcal{P},\mathcal{A}})$ for the skills $\Omega$.
After a period of continual learning in one or more settings, a robot may be evaluated on a task-level planning problem $\mathbf{p}=({{s}_{0},\mathcal{O},\mathcal{S}_{g}})$, where $\mathcal{S}_{g}\subseteq\mathcal{S}$ is the set of goal states to be reached. Given a model $\mathcal{M}=({\mathcal{P},\mathcal{A}})$, a classical planner can be used to search for an abstract plan $\underline{a}_{1},\dots,\underline{a}_{n}$}] that solves the abstract planning problem $\underline{\mathbf{p}}=({{\underline{s}}_{0},\mathcal{O},\underline{\mathcal{S}}_{g}})$.
Definition 2.
An abstract plan $\underline{a}_{1},\dots,\underline{a}_{n}$}] is a solution for skill planning problem $\mathbf{p}=({{s}_{0},\mathcal{O},\mathcal{S}_{g}})$ iff for $1\leq i\leq n,\,s_{i}=T(s_{i-1},\underline{\omega}_{i}),s_{i-1}\in\mathcal{I}_{\underline{\omega}_{i}},\text{ and }s_{n}\in\mathcal{S}_{g},$ where $\underline{\omega}_{i}=\underline{\omega}_{\underline{a}_{i}}$.
In Sec. III, we describe how SkillWrapper constructs $\mathcal{M}$} from raw skill executions.
III SkillWrapper
@@
In this section, we introduce SkillWrapper, a novel approach that autonomously learns symbolic representations for black-box skills while providing several guarantees such as soundness and completeness (see Appendix -B for further theoretical details). To produce a valid abstract model that enables planning,  SkillWrapper iterates through a three-step process: 1) actively proposing and executing exploratory skill sequences to collect data on the initiation and termination set of each skill, 2) incrementally building a set of predicates from scratch by contrasting positive and negative examples, and then 3) constructing valid operators using these invented predicates, from which further exploratory skill sequences can be proposed. This procedure is outlined in Algorithm 1. As SkillWrapper continues to collect data, invent and incorporate new predicates, and update its planning model, it progressively improves the accuracy of the learned task-level transition model. The resulting task-level representations, or operators, can be used with an off-the-shelf classical planner to solve task planning problems specified using RGB images. We delve into the core components of our system (lines 4–9 of Algorithm 1) in the following subsections. Importantly, we provide strong theoretical results for the soundness and completeness of SkillWrapper in Section III-D.
@@
9:return $\mathcal{M}$
III-A Active Data Collection
To collect data for learning, our method computes a sequence of skills $\langle\underline{\omega}_{1},\dots,\underline{\omega}_{k}\rangle$ for the robot to execute and collects the resulting transitions. While executing these skill sequences, the robot collects a dataset $\mathcal{D}$ of transitions of the form $\langle s,\underline{\omega},s^{\prime}\rangle$, where $s,s^{\prime}\in{\mathcal{S}}$. These transitions can answer two questions: